OpenStudy (cam1258):

What value of k makes the equation true? (5a²b³)(6a^k(b))=30a^6(b^4) A)2 B)3 C)4 D)8

jhonyy9 (jhonyy9):

there are (5a^2b^3)(6a^(k) *b) = 30a^6 *b^4 30a^(2k) *b^4 = 30a^6 *b^4 2k = 6 k = ?

jhonyy9 (jhonyy9):

do you understand me ?

jhonyy9 (jhonyy9):

@cam1258 do you have question about above wrote ?

OpenStudy (cam1258):

I know it is not A or D

OpenStudy (cam1258):

Is that correct?

OpenStudy (seratul):

If 2k=6, k=3...

jhonyy9 (jhonyy9):

so than you see there above wrote that 2k = 6 divide both sides by 2 k = 6/2 = ?

OpenStudy (cam1258):

B

jhonyy9 (jhonyy9):

@cam1258 do you understand the way how you get the right answer ?

OpenStudy (cam1258):

Yes, I do. Thank you for explaining and helping

jhonyy9 (jhonyy9):

yw my pleasure and anytime good luck bye bye

OpenStudy (mathstudent55):

\((5a^2b^3)(6a^k(b))=30a^6(b^4)\) \(5 \times a^2 \times b^3 \times 6 \times a^k \times b=30 \times a^6 \times b^4\) \(5 \times 6 \times a^2 \times a^k \times b^3 \times b=30 \times a^6 \times b^4\) \((5 \times 6) \times (a^2 \times a^k) \times (b^3 \times b)=30 \times a^6 \times b^4\) \(30 \times a^{2+k} \times b^4 = 30 \times a^6 \times b^4\) \(\cancel{30} \times a^{2+k} \times \cancel{b^4} = \cancel{30} \times a^6 \times \cancel{b^4}\) \(a^{2+k} = a^6\) \(2 + k = 6\) \(k = 4\)

OpenStudy (mathstudent55):

@jhonyy9 \(\Large a^2 \times a^k = a^{2 + k} \ne a^{2k} \)

OpenStudy (mathstudent55):

Answer is not B.

jhonyy9 (jhonyy9):

ohhh - yes ! sorry - but the way is to start like above i started it so this mean that a^(2+k) = a^6 so 2+k=6 => k = 6-2 so k=4 and in this way the choice C. is right sure thank you very very much @mathstudent55